How much salicylic acid (the limiting reagent in the synthesis) is need to produce 500 g of aspirin?

Mw aspirin is 180.158 g/mol

Mw salicylic acid is 138.12 g/mol

a.) 38.333

b.) 383.33

c.) 3.833

d.) 0.3833
Answer #1

C7H6O3 (salicylic acid) + C4H6O3 = C9H8O4 (aspirin) + HC2H3O2

In the above equation, one mole of salicylic acid gives one mole of aspirin.

180.158 g of aspirin = 138.12 g of salicylic acid

1 g aspirin = 138.12/180.158 g salicylic acid

500 g aspirin = (138.12/180.158) × 500 =383.33 g salicylic acid

Option B is correct

Please rate.